How many grams of N2 can be produce when 4.00 g of O2 how to react
1
Expert's answer
2020-04-20T15:08:30-0400

In order to understand how much N2 can be produce, the reaction equation must be specified. Consider the combustion of ammonia:


4NH3 + 3O2 = 2N2 + 6H2O


If react 4.00/32 = 0.125 mol we get 0.125*3/2 = 0.1875 mol of N2. Mass of N2 is 0.1875*28 = 5.25 g.